OpenStudy (anonymous):

solve graphically y = (1/2)^(-x-2) = -x^2 + 5

OpenStudy (susanna):

\[y=(\frac{ 1 }{ 2 })^{-x-2}=-x ^{2}+5\]

OpenStudy (tkhunny):

Stat with this: \(\left(\dfrac{1}{2}\right)^{-x-2} = \left(\dfrac{1}{2}\right)^{-(x+2)} = \left(\left(\dfrac{1}{2}\right)^{-1}\right)^{(x+2)} = 2^{x+2}\).

OpenStudy (anonymous):

alright, did that, what is next?

OpenStudy (triciaal):

plot 2 graphs on the same grid y = 1/2^(-x-2) and y = -x^2 + 5 use the same set of x values find where the graphs intersect

OpenStudy (anonymous):

thank you again triciaal and tkhunny

TheSmartOne (thesmartone):

https://www.desmos.com/calculator/ejeeqrztmz

OpenStudy (triciaal):

you are welcome!!